OpenStudy (anonymous):

x + 9 = 18 - 2x Solution: x = 3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

What are you doing here? Please explain (:

myininaya (myininaya):

want us to check it? 3+9=18-2(3) 12=18-6 12=12 that is correct

OpenStudy (gw2011):

x+9=18-2x x+2x+9-18=0 3x-9=0 3x=9 x=3

OpenStudy (anonymous):

yeah, what is the question?